OpenStudy (anonymous):

If A can do a job in 8 days and b can do the same job in 12 days, how long would it take them both to complete the job working together?

OpenStudy (anonymous):

8 + 12 = 20 would take 20 days. \[\color{red}{\text{ok joemath, take it away}}\]

OpenStudy (anonymous):

\[\frac{1}{\frac{1}{8}+\frac{1}{12}}\]

OpenStudy (anonymous):

just kidding. solve \[(\frac{1}{8}+\frac{1}{12})t=1\] for t

OpenStudy (anonymous):

one day i am going to figure out a snap way to explain this.

OpenStudy (anonymous):

ends up being \[3x+4x=24\] but why this is obvious eludes me. joemath?

OpenStudy (anonymous):

yeah >.< i just think of it like: d = rt the rate is inversely proportional to time, so i need to get the reciprocals of their times to get the "rates". Then i add the rates together, but i want time as an answer, so i take the reciprocal of my answer...... its gross.

OpenStudy (anonymous):

you should have seem me with the jellybean problem. took maybe half an hour!

OpenStudy (anonymous):

sounds intense o.O

OpenStudy (anonymous):

no joemath no one thinks like that. must be some obvious way to do it. i mean look at the answer. right hand side of equal sign in lcm of 8 and 12. left hand side is \[\frac{24}{8}+\frac{24}{12}\] which for one thing means my solution is wrong. but still there must be some obvious reason for this yes?

OpenStudy (anonymous):

i guess it has to do with the fact that you would use the LCM to get a common denominator for the fractions? and by the way, I think like that :( lol